# Solitons and singularities in relativistic ultrastiff fluids

### v1

@@ I Introduction
There is a growing consensus that the speed of sound ($c_{s}$) of neutron-star matter is quite high, potentially exceeding 0.6 [1, 2] (with $c=1$), though the extent of this excess remains uncertain. This finding strengthens the widespread belief that nothing should prevent the speed of sound of dense matter from approaching the speed of light in some regions of the QCD phase diagram [3, 4, 5] [^1] A novel upper bound (namely $c_{s}^{2}\leq 0.781$) has been recently introduced in [6]. However, that is not a statement about the non-existence of fluids with larger speed of sound. It just provides a dynamical characterization of such fluids. In particular: Fluids with $c_{s}^{2}>0.781$ cannot be modeled using ordinary Israel-Stewart theory [7, 8], or any kind of viscoelastic Maxwell-type model [9, 10]. Currently, the only universally accepted existence bound is $c_{s}^{2}\leq 1$, which can be derived directly from relativistic statistical mechanics [11, 12, 13].. This is a rather intriguing possibility since, in a phase of this kind, sound waves would propagate with the same speed in all reference frames, potentially giving rise to unusual fluid-dynamical phenomena that are intrinsically relativistic. The goal of this article is to analyse some of these phenomena.
Most instances of fluid phases with luminal sound speed discussed in the literature [3, 4, 5] may be approximated as “ultrastiff fluids”, namely fluids whose equation is approximately given by
@@ II.2 Thermodynamic interpretation
The first equation tells us that the baryon number is conserved, while the second tells us that the “chemical momentum” is the gradient of a scalar, which is just a restatement of the irrotationality assumption. The latter condition is known to occur in zero-temperature superfluids [21, 22, 23, 24] (note that, at $T=0$, the normal component disappears), where $\Psi$ is proportional to the gradient of the phase of the order parameter. Then, the dual scalar-field theory is just the effective field theory of the massless Goldstone boson associated with the superfluid symmetry breaking. However, we remark that superfluidity is only a sufficient condition for irrationality. It is not a necessary condition. Any zero-temperature ideal fluid whose initial state happens to be irrotational remains irrotational at all times.
@@ IV.2 Solitonic behavior
Figure 2: Non-linear interaction of two solitonic waves (blue), compared to the naive prediction using linearised hydrodynamics (dashed). The background state has pressure $A^{2}$, the right-moving soliton has profile $f(L)=Ae^{-L^{4}}$, and right-moving soliton has profile $g(L)=Ae^{-4L^{3/2}}$. When the solitons meet, non-linearities cause the pressure to become much larger than the sum of the pressures of the individual waves. Despite this non-linearity, the waves recover their original shape after the interaction.

### v2

@@ I Introduction
There is a growing consensus that the speed of sound ($c_{s}$) of neutron-star matter is quite high, potentially exceeding 0.6 [1, 2] (with $c=1$), though the extent of this excess remains uncertain. This finding strengthens the widespread belief that nothing prevents the speed of sound of dense matter from approaching the speed of light in some regions of the QCD phase diagram [3, 4, 5] [^1] A novel upper bound (namely $c_{s}^{2}\leq 0.781$) has been recently introduced in [6]. However, that is not a statement about the non-existence of fluids with larger speed of sound. It just provides a dynamical characterization of such fluids. In particular: Fluids with $c_{s}^{2}>0.781$ cannot be modeled using ordinary Israel-Stewart theory [7, 8], or any kind of viscoelastic Maxwell-type model [9, 10]. Currently, the only universally accepted existence bound is $c_{s}^{2}\leq 1$, which can be derived directly from relativistic statistical mechanics [11, 12, 13].. This is a rather intriguing possibility since, in a phase of this kind, sound waves would propagate with the same speed in all reference frames, potentially giving rise to unusual fluid-dynamical phenomena that are intrinsically relativistic. The goal of this article is to analyse some of these phenomena.
Most instances of fluid phases with luminal sound speed discussed in the literature [3, 4, 5] may be approximated as “ultrastiff fluids”, namely fluids whose equation of state is approximately given by
@@ II.2 Thermodynamic interpretation
The first equation tells us that the baryon number is conserved, while the second tells us that the “chemical momentum” is the gradient of a scalar, which is just a restatement of the irrotationality assumption. The latter condition is known to occur in zero-temperature superfluids [21, 22, 23, 24] (note that, at $T=0$, the normal component disappears), where $\Psi$ is proportional to the gradient of the phase of the order parameter. Then, the dual scalar-field theory is just the effective field theory of the massless Goldstone boson associated with the superfluid symmetry breaking. However, we remark that superfluidity is only a sufficient condition for irrotationality. It is not a necessary condition. Any zero-temperature ideal fluid whose initial state happens to be irrotational remains irrotational at all times.
@@ IV.2 Solitonic behavior
Figure 2: Non-linear interaction of two solitonic waves (blue), compared to the naive prediction using linearised hydrodynamics (dashed). The background state has pressure $A^{2}$, the right-moving soliton has profile $f(L)=Ae^{-L^{4}}$, and left-moving soliton has profile $g(L)=Ae^{-4L^{3/2}}$. When the solitons meet, non-linearities cause the pressure to become much larger than the sum of the pressures of the individual waves. Despite this non-linearity, the waves recover their original shape after the interaction.